A cake shop has cookies worth $9.00 per pound and cookies worth $15.00 per pound. How many pounds of each type of cookie should the manager mix together to make 120 pounds of cookies at $12.50 per Pound?
Answer by stanbon(50474)

A cake shop has cookies worth $9.00 per pound and cookies worth $15.00 per pound. How many pounds of each type of cookie should the manager mix together to make 120 pounds of cookies at $12.50 per Pound?
-----
Quantity Eq: n + f = 120 lbs
Value Eq:: 9n+15f = 12.5*120
------------------------------------
Multiply thru Quantity by 9 to get:
9n + 9f = 9*120
9n +15f = 12.5*120
------
6f = 3.5*120
f = 20*3.5
f = 70 (# of lbs. of $15 cookies)
Solve for "n":
n + f = 120
n = 120-70 = 50 lb (# of lbs. of $9 cookies)
=====================
Cheers,
San H.